# SVGDreamer: Text Guided SVG Generation with Diffusion Model

### v4

@@ Abstract
Recently, text-guided scalable vector graphics (SVGs) synthesis has shown promise in domains such as iconography and sketch. However, existing text-to-SVG generation methods lack editability and struggle with visual quality and result diversity. To address these limitations, we propose a novel text-guided vector graphics synthesis method called SVGDreamer. SVGDreamer incorporates a semantic-driven image vectorization (SIVE) process that enables the decomposition of synthesis into foreground objects and background, thereby enhancing editability. Specifically, the SIVE process introduce attention-based primitive control and an attention-mask loss function for effective control and manipulation of individual elements. Additionally, we propose a Vectorized Particle-based Score Distillation (VPSD) approach to tackle the challenges of shape over-smoothing, color over-saturation, limited diversity in results, and slow convergence in existing text-to-SVG generation methods. VPSD models SVGs as distributions of control points and colors to counteract over-smoothing and over-saturation. Furthermore, VPSD leverages a reward model to reweight vector particles, which improves aesthetic appeal and accelerates convergence. Extensive experiments have been conducted to validate the effectiveness of SVGDreamer, demonstrating its superiority over baseline methods in terms of editability, visual quality, and diversity. Project page: https://ximinng.github.io/SVGDreamer-project/
@@ 1 Introduction
Despite their impressive performance, existing T2I-based methods have certain limitations. Firstly, the vector images generated by these methods lack editability. Unlike the conventional approach of creating vector graphics, where individual elements are added one by one, T2I-based methods do not distinguish between different components during synthesis. As a result, the objects become entangled, making it challenging to edit or modify a single object independently. Secondly, there is still a large room for improvement in visual quality and diversity of the results generated by these methods. Both VectorFusion [12] and DiffSketcher [48] use data augmentation extended Score Distillation Sampling (SDS) [23] to distill priors from the T2I models. However, it has been observed that SDS can lead to issues such as color over-saturation and over-smoothing, resulting in a lack of fine details in the generated vector images. Besides, SDS optimizes a set of control points in the vector graphic space to obtain the average state of the vector graphic corresponding to the text prompt in a mode-seeking manner [23]. This leads to a lack of diversity and detailed construction in the SDS-based approach [12, 48], along with absent text prompt objects.
@@
To improve the visual quality and diversity of the generated vector graphics, we introduce a new technique: Vectorized Particle-based Score Distillation (VPSD) for vector graphics refinement. Previous works in vector graphics synthesis [12, 48, 11] that utilized score distillation sampling [23] often encountered issues like shape over-smoothing, color over-saturation, limited diversity in results, and slow convergence in the synthesis results [23, 48]. To address these challenges, VPSD models SVGs as distributions of control points and colors, respectively. VPSD adopts a LoRA [10] network to estimate the distributions, aligning vector graphics with the pretrained diffusion model. Furthermore, to enhance the aesthetic appeal of the generated vector graphics, we introduce ReFL [49] to fine-tune the estimation network. Through this refinement process, we achieve the final vector graphics that exhibit high editability, superior visual quality, and increased diversity. To validate the effectiveness of our proposed method, we perform extensive experiments to evaluate the model across multiple aspects. In summary, our contributions can be summarized as follows:
@@ 3 Methodology
Figure 2: Overview of SVGDreamer. The method consists of two parts: semantic-driven image vectorization (SIVE) and SVG synthesis through VPSD optimization. The result obtained from SIVE can be used as input for VPSD, which helps in establishing a dependable initialization point.
@@ 3.2 Vectorized Particle-based Score Distillation
Figure 3: An illustration of the VPSD. VPSD accepts $k$ sets of SVG parameters as input. The rendered image is sent to the pretrained diffusion and the score of the SVG parameters distribution to compute the gradient of VPSD.
@@
In [45], only selected particles update the LoRA network in each iteration. However, this approach neglects the learning progression of LoRA networks, which are used to represent variational distributions. These networks typically require numerous iterations to approximate the optimal distribution, resulting in slow convergence. Unfortunately, the randomness introduced by particle initialization can lead to early learning of suboptimal particles, which adversely affects the final convergence result. In VPSD, we introduce a Reward Feedback Learning method, as Fig. 3 illustrates. This method the leverages a pre-trained reward model [49] to assign reward scores to samples collected from LoRA model. Then LoRA model subsequently updates from these reweighted samples,
@@
where $g_{\phi_{\mathrm{est}}}(y)$ denotes the generated image of $\mu$ model with parameters $\phi_{\mathrm{est}}$ corresponding to prompt $y$, and $r$ represents the pretrained reward model [49], $\psi$ represents reward-to-loss map function implemented by ReLU, and $\lambda=1e-3$. We used the DPM-solver [17] to rapidly sample $k$ samples during the early iteration stage. This method saves 2 times the iteration step for VPSD convergence and improves the aesthetic score of the SVG by filtering out samples with low reward values in LoRA.
@@
Figure 4: Results of qualitative evaluation. Note: DiffSketcher is originally designed for vector sketch generation. We reimplemented it for generating RGB vector images.
@@ 4.1 Qualitative Evaluation
Table 1: Results of the quantitative evaluation.
Method / Metric FID [6$\downarrow$ PSNR [9$\uparrow$ CLIPScore [24$\uparrow$ BLIPScore [13$\uparrow$ Aesthetic [30$\uparrow$ HPS [47$\uparrow$
@@
Figure 5: Examples of vector assets created by our SVGDreamer model.
Figure 4 presents a qualitative comparison between SVGDreamer and existing text-to-SVG methods. Compared to CLIPDraw [4], SVGDreamer synthesizes SVGs with higher fidelity and detail. And we compare our work with SDS-based methods [12, 48], emphasizing our ability to address challenges such as excessive smoothing and color over-saturation. As shown in the 5th column, SIVE achieves semantic decoupling, but it is unable to overcome the smooth nature of SDS. As observed in the last two columns, our approach demonstrates superior detail compared to the SDS-based approach, regardless of whether it was optimized from scratch or through the full process. Consequently, this results in a higher aesthetic score.
@@ 4.2 Quantitative Evaluation
To demonstrate the effectiveness of our proposed method, we conduct comprehensive experiments to evaluate the model across various aspects, including Fréchet Inception Distance (FID) [6], Peak Signal-to-Noise Ratio (PSNR) [9], CLIPScore [24], BLIPScore [13], Aesthetic score [30] and Human Performance Score [47]. Table 1 presents a comparison of our approach with the most representative text-to-SVG methods, including CLIPDraw [4], VectorFusion [12], and DiffSketcher [48]. We conducted a quantitative evaluation of the six styles identified in Sec. 3.3, with each style comprising 10 unique prompts and 50 synthesized SVGs per prompt. For diversity evaluation of vector graphics and fill color saturation, we used SD sampling results as a Ground Truth (GT) and calculated FID and PSNR metrics respectively. The quantitative analysis in the first two columns indicates that our method surpasses other methods in terms of FID and PSNR. This suggests that our method offers a greater range of diversity compared to SDS-based synthesis [12, 48]. To assess the consistency between the generated SVGs and the provided text prompts, we used both CLIPScore and BLIPScore. To measure the perceptual quality of synthetic vector images, we measure aesthetic scores using the LAION aesthetic classifier [30]. In addition, we introduced human performance score [30] (HPS) to evaluate our approach from a human aesthetic perspective.
@@ 4.3 Ablation Study
Figure 6: Comparison of LIVE vectorization with SIVE. In the first row, “Foreground 1 (or 2)” indicates Astronaut and Plants, respectively. Glyphs are added manually rather than produced by our method. In LIVE setup, we follow the LIVE protocol in VectorFusion [12], representing a vector image with 128 paths distributed across four layers of 32 paths each.
@@
4.3.2 VPSD v.s. LSDS [12, 11] v.s ASDS [48]
@@ 5 Conclusion
In this work, we have introduced SVGDreamer, an innovative model for text-guided vector graphics synthesis. SVGDreamer incorporates two crucial technical designs: semantic-driven image vectorization (SIVE) and vectorized particle-based score distillation (VPSD), which empower our model to generate vector graphics with high editability, superior visual quality, and notable diversity. SVGDreamer is expected to significantly advance the applications of text-to-SVG models in the design field.
Limitations. Our method’s editability, dependent on the T2I model used, is currently limited. However, future advancements in T2I diffusion models could improve our model’s decomposition abilities, extending its editability. Furthermore, studying how to automatically determine the number of control points at the SIVE object level is valuable. We believe this will contribute to the advancement of semantically guided image vectorization.
6 Acknowledgement
This work is supported by the CCF-Baidu Open Fund Project and Young Elite Scientists Sponsorship Program by CAST.
@@ 6 Acknowledgement
  • In section D, We provide a guide on how to identify semantic objects in text prompts.
@@
  • In section F, we provide the experiments of VPSD in raster image synthesis.
@@ Appendix B Applications of SVGDreamer
Figure 8: Additional qualitative results of SVGDreamer.
@@ Appendix C Implementation Details
Appendix D Object Identification in SIVE Prompt
@@
It is common for multiple nouns within a sentence to refer to the same object. In our experiments, we did not employ a specific selection strategy because the cross-attention maps for such nouns-for example, “man” and “astronaut” – are very similar. We present two examples in Fig. 11. Therefore, choosing either “man” or “astronaut” produces similar results with our method. For more precise control, users may utilize the cross-attention maps of the text prompt to identify the desired objects. In SIVE, users can use visual text prompts to identify semantic objects.
Appendix E Additional Ablation Study
@@ E.2 Ablation on ReFL
Figure 13: Effects of the Reward Feedback Learning (ReFL) method on the generated results. When employing the ReFL method, the visual quality of the generated results is significantly enhanced.
In [45], only selected particles update the LoRA network in each iteration. However, this approach neglects the learning progression of LoRA networks, which are used to represent variational distributions. These networks typically require numerous iterations to approximate the optimal distribution, resulting in slow convergence. Unfortunately, the randomness introduced by particle initialization can lead to early learning of sub-optimal particles, which adversely affects the final convergence result. In VPSD, we introduce a Reward Feedback Learning (ReFL) method. This method leverages a pre-trained reward model [49] to assign reward scores to samples collected from LoRA model. Then LoRA model subsequently updates from these reweighted samples. As indicated in Table 2, this led to a significant reduction in the number of iterations by almost 50%, resulting in a 50% decrease in optimization time. And improves the aesthetic score of the SVG by filtering out samples with low reward values in LoRA. Filtering out samples with low reward values, as demonstrated in Tab. 1, enhances the aesthetic score of the SVG. The visual improvements brought by ReFL are illustrated in Fig. 13.
@@
E.3 Ablation on Number of Vector Particles
@@ E.4 Ablation on the Number of Paths
Figure 15: The number of paths can influence the style of the vector graphics.
Figure 16: Comparison of the results of VSD and VPSD in image synthesis.
Appendix F VPSD in 2D image synthesis
VPSD is specifically designed for text-to-SVG tasks; however, it can also be adapted for 2D image synthesis. As illustrated in Fig. 16, images synthesized by VSD may exhibit displaced or incomplete object layouts, resulting in samples that might not meet human aesthetic preferences. In contrast, VPSD integrates a reward score within its feedback learning process, which significantly enhances the quality of the generated images.

### v5

@@ Abstract
Recently, text-guided scalable vector graphics (SVGs) synthesis has shown promise in domains such as iconography and sketch. However, existing text-to-SVG generation methods lack editability and struggle with visual quality and result diversity. To address these limitations, we propose a novel text-guided vector graphics synthesis method called SVGDreamer. SVGDreamer incorporates a semantic-driven image vectorization (SIVE) process that enables the decomposition of synthesis into foreground objects and background, thereby enhancing editability. Specifically, the SIVE process introduces attention-based primitive control and an attention-mask loss function for effective control and manipulation of individual elements. Additionally, we propose a Vectorized Particle-based Score Distillation (VPSD) approach to address issues of shape over-smoothing, color over-saturation, limited diversity, and slow convergence of the existing text-to-SVG generation methods by modeling SVGs as distributions of control points and colors. Furthermore, VPSD leverages a reward model to re-weight vector particles, which improves aesthetic appeal and accelerates convergence. Extensive experiments are conducted to validate the effectiveness of SVGDreamer, demonstrating its superiority over baseline methods in terms of editability, visual quality, and diversity. Project page: https://ximinng.github.io/SVGDreamer-project/
@@ 1 Introduction
Despite their impressive performance, existing T2I-based methods have certain limitations. Firstly, the vector images generated by these methods lack editability. Unlike the conventional approach of creating vector graphics, where individual elements are added one by one, T2I-based methods do not distinguish between different components during synthesis. As a result, the objects become entangled, making it challenging to edit or modify a single object independently. Secondly, there is still a large room for improvement in visual quality and diversity of the results generated by these methods. Both VectorFusion [12] and DiffSketcher [48] extended the Score Distillation Sampling (SDS) [22] to distill priors from the T2I models. However, it has been observed that SDS can lead to issues such as color over-saturation and over-smoothing, resulting in a lack of fine details in the generated vector images. Besides, SDS optimizes a set of control points in the vector graphic space to obtain the average state of the vector graphic corresponding to the text prompt in a mode-seeking manner [22]. This leads to a lack of diversity and detailed construction in the SDS-based approach [12, 48], along with absent text prompt objects.
@@
To improve the visual quality and diversity of the generated vector graphics, we introduce Vectorized Particle-based Score Distillation (VPSD) for vector graphics refinement. Previous works in vector graphics synthesis [12, 48, 11] that utilized SDS often encountered issues like shape over-smoothing, color over-saturation, limited diversity, and slow convergence in synthesized results [22, 48]. To address these issues, VPSD models SVGs as distributions of control points and colors, respectively. VPSD adopts a LoRA [10] network to estimate these distributions, aligning vector graphics with the pretrained diffusion model. Furthermore, to enhance the aesthetic appeal of the generated vector graphics, we integrate ReFL [49] to fine-tune the estimation network. Through this refinement process, we achieve final vector graphics that exhibit high editability, superior visual quality, and increased diversity. To validate the effectiveness of our proposed method, we perform extensive experiments to evaluate the model across multiple aspects. In summary, our contributions can be summarized as follows:
@@ 3 Methodology
Figure 2: Overview of SVGDreamer. The method consists of two parts: semantic-driven image vectorization (SIVE) and SVG synthesis through VPSD optimization. The result obtained from SIVE can be used as input of VPSD for further refinement.
@@ 3.2 Vectorized Particle-based Score Distillation
Figure 3: The process of VPSD.
@@
In [45], only randomly selected particles update the LoRA network in each iteration. However, this approach neglects the learning progression of vector particles, which are used to represent the optimal SVG distributions. Furthermore, these networks typically require numerous iterations to approximate the theoretical optimal distribution, resulting in slow convergence. In VPSD, we introduce a Reward Feedback Learning method, as Fig. 3 illustrates. This method leverages a pre-trained reward model [49] to assign reward scores to samples collected from LoRA model. Then LoRA model subsequently updates from these reweighted samples,
@@
where $g_{\phi_{\mathrm{est}}}(y)$ denotes the generated image of $\mu$ model with parameters $\phi_{\mathrm{est}}$ corresponding to prompt $y$, and $r$ represents the pretrained reward model [49], $\psi$ represents reward-to-loss map function implemented by ReLU, and $\lambda=1e-3$. We used the DDIM [32] to rapidly sample $k$ samples during the early iteration stage. This method saves 2 times the iteration step for VPSD convergence and improves the aesthetic score of the SVG by filtering out samples with low reward values in LoRA.
@@
Figure 4: Qualitative comparison of different methods. Note that DiffSketcher was originally designed for vector sketch generation; therefore, we re-implemented it to generate RGB vector images.
@@ 4.1 Qualitative Evaluation
Table 1: Quantitative comparison of different methods.
Method / Metric FID [6$\downarrow$ PSNR [9$\uparrow$ CLIPScore [23$\uparrow$ BLIPScore [13$\uparrow$ Aesthetic [29$\uparrow$ HPS [47$\uparrow$
@@
Figure 5: Examples of vector assets created by our SVGDreamer.
Figure 4 presents a qualitative comparison between SVGDreamer and existing text-to-SVG methods. Compared to CLIPDraw [4], SVGDreamer synthesizes SVGs with higher fidelity and detail. We also compare our work with SDS-based methods [12, 48], emphasizing our ability to address issues such as shape over-smoothing and color over-saturation. As shown in the fifth column, SIVE achieves semantic decoupling but cannot overcome the inherently smooth nature of SDS. As observed in the last two columns, our approach demonstrates superior detail compared to the SDS-based approach, regardless of whether the model was optimized from scratch or through the entire process. Consequently, this leads to a higher aesthetic score.
@@ 4.2 Quantitative Evaluation
To demonstrate the effectiveness of our proposed method, we conducted comprehensive experiments to evaluate the model across various aspects, including Fréchet Inception Distance (FID) [6], Peak Signal-to-Noise Ratio (PSNR) [9], CLIPScore [23], BLIPScore [13], Aesthetic score [29] and Human Performance Score [47] (HPS). Table 1 presents a comparison of our approach with the most representative text-to-SVG methods, including CLIPDraw [4], VectorFusion [12], and DiffSketcher [48]. We conducted a quantitative evaluation of the six styles identified in Sec. 3.3, with each style comprising 10 unique prompts and 50 synthesized SVGs per prompt. For diversity evaluation of vector graphics and fill color saturation, we used SD sampling results as a Ground Truth (GT) and calculated FID and PSNR metrics respectively. The quantitative analysis in the first two columns indicates that our method surpasses other methods in terms of FID and PSNR. This suggests that our method offers a greater range of diversity compared to SDS-based synthesis [12, 48]. To assess the consistency between the generated SVGs and the provided text prompts, we used both CLIPScore and BLIPScore. To measure the perceptual quality of synthetic vector images, we measure aesthetic scores using the LAION aesthetic classifier [29]. Besides, we use HPS to evaluate our approach from a human aesthetic perspective.
@@ 4.3 Ablation Study
Figure 6: Comparison of LIVE vectorization with SIVE. In the first row, ”Foreground 1” and ”Foreground 2” refer to Astronaut and Plants, respectively. Glyphs have been added manually and were not produced by our method. In the LIVE setup, we follow the protocol outlined in VectorFusion [12], which represents a vector image with 128 paths distributed across four layers, with 32 paths in each layer.
@@
4.3.2 VPSD v.s. LSDS [12, 11] v.s. ASDS [48]
@@ 5 Conclusion
In this work, we have introduced SVGDreamer, an innovative model for text-guided vector graphics synthesis. SVGDreamer incorporates two crucial technical designs: Semantic-Driven Image Vectorization (SIVE) and Vectorized Particle-Based Score Distillation (VPSD). These empower our model to generate vector graphics with high editability, superior visual quality, and notable diversity. SVGDreamer is expected to significantly advance the application of text-to-SVG models in the design field.
Limitations. The editability of our method, which depends on the text-to-image (T2I) model used, is currently limited. However, future advancements in T2I diffusion models could enhance the decomposition capabilities of our approach, thereby extending its editability. Moreover, exploring ways to automatically determine the number of control points at the SIVE object level is valuable.
Acknowledgement. This work is supported by the CCF-Baidu Open Fund Project and Young Elite Scientists Sponsorship Program by CAST.
@@
  • In section D, We explain how to identify semantic objects in SIVE prompts.
@@
  • In section F, we provide example results from using VPSD for raster image synthesis.
@@ Appendix B Applications of SVGDreamer
Figure 8: More results generated by our SVGDreamer.
@@ Appendix C Implementation Details
Appendix D Object Identification in SIVE Prompts
@@
It is common for multiple nouns within a sentence to refer to the same object. We present two examples in Fig. 11. In our experiments, we did not employ a specific selection strategy because the cross-attention maps for such nouns-for example, “man” and “astronaut” – are very similar. Therefore, choosing either “man” or “astronaut” produces similar results with our method. For more precise control, users may utilize the cross-attention maps of the text prompt to identify the desired objects. In SIVE, users can use visual text prompts to identify semantic objects.
Appendix E Additional Ablation Studies
@@ E.2 Ablation on ReFL
Figure 13: Effect of the Reward Feedback Learning (ReFL) on the generated results. When employing ReFL, the visual quality of the generated results is significantly enhanced.
In [45], only selected particles update the LoRA network in each iteration. However, this approach neglects the learning progression of LoRA networks, which are used to represent variational distributions. These networks typically require numerous iterations to approximate the optimal distribution, resulting in slow convergence. Unfortunately, the randomness introduced by particle initialization can lead to early learning of sub-optimal particles, which adversely affects the final convergence result. In VPSD, we introduce a Reward Feedback Learning (ReFL) method. This method leverages a pre-trained reward model [49] to assign reward scores to samples collected from LoRA model. Then LoRA model subsequently updates from these reweighted samples. As indicated in Table 2, this led to a significant reduction in the number of iterations by almost 50%, resulting in a 50% decrease in optimization time. And improves the aesthetic score of the SVG by filtering out samples with low reward values in LoRA. Filtering out samples with low reward values, as demonstrated in Table 1, enhances the aesthetic score of the SVG. The visual improvements brought by ReFL are illustrated in Fig. 13.
@@
E.3 Ablation on the Number of Vector Particles
@@ E.4 Ablation on the Number of Paths
Figure 15: Effect of the number of paths on synthesized results.
Figure 16: Comparison of the results from using VPSD and VSD for 2D image synthesis.
Appendix F VPSD for 2D Image Synthesis
In this work, VPSD is specifically designed for text-to-SVG generation; however, it can also be adapted for 2D image synthesis. As illustrated in Fig. 16, images synthesized by VSD may exhibit displaced or incomplete object layouts, resulting in samples that might not meet human aesthetic preferences. In contrast, VPSD integrates a reward score within its feedback learning process, which significantly enhances the quality of the generated images.